# Sharp Self-Normalized Concentration Inequalities of Marginal Mean with Sample Variance Only

### v2

@@ Abstract
(This is the first version of a working paper. A more detailed follow-up with applications and uniform results is in preparation.) We develop a family of self-normalized concentration inequalities for marginal mean under martingale-difference structure and $\phi/\tilde{\phi}-$mixing conditions, where the latter includes many processes that are not strongly mixing. The variance term is fully data-observable: naive sample variance in the martingale case and an empirical block long-run variance under mixing conditions. Thus, no predictable variance proxy is required. No specific assumption on the decay of the mixing coefficients (e.g. summability) is needed for the validity. The constants are explicit and the bounds are ready to use.
@@ 1 Literature Review
This section is dedicated to the literature review of the previous research on self-normalized and empirical Bernstein (EB) inequalities for marginal mean of real-valued random variables. Marginal mean here means $\mathbb{E}[Z_{1}]$ (or $\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}[Z_{i}]$) when samples $Z_{i}$’s have common (or different) mean(s).
@@ 1.1 Self-Normalized Inequality
They also proved that (3) holds for “heavy-on-left” MDS as well. Compared with (1), (2)-(4) hold under much weaker moment conditions but their statistical applications are not so convenient as (1). the dependence on $\mathbf{M}]_{n}$ appears inside the Laplace transform or in the event $\{[\mathbf{M}]_{n}\leq y\}$, and there is no closed-form expression of the type $\sqrt{[\mathbf{M}]_{n}\log(\frac{1}{\delta})}$ with an explicit confidence parameter $\delta$. However, the variance ($\sigma$) and boundedness ($b$) are usually unknown to us, which reject the direct application of (1).
@@ 1.2 Empirical Bernstein Inequality of Marginal Mean
$\sqrt{n}L_{n}\xrightarrow[]{\mathbb{P}}\sqrt{2\log(1/\alpha)\sigma^{2}},$ (5)
@@
In summary, the results of Ramdas and co-authors are indeed elegant. But, as we discussed for fixed-n EB inequality of marginal mean, they have three structural limitations: (i) they require a known and tight bound on the support length to calculate the confidence interval; (ii) they primarily target the predictable mean (the arithmetic average of conditional expectations), and in general mixing settings there is no clear high-probability control of the gap between this predictable mean and the marginal mean—unless one assumes that all conditional expectations are constant, which essentially reduces to the martingale-difference case. Their recent important work, e.g. ? (?) and ? (?), still rely on the assumption of constant conditional expectation. (iii) they rely on a carefully tuned variance proxy satisfying the following three points simultaneously, (i) predictable; (ii) ensure the super-martingale property; (iii) converge to the true variance. The design of this proxy is highly non-trivial even for martingale differences or i.i.d. data. For some complex mixing processes, whose relevant variance of partial sum is the long-run variance rather than the marginal variance, the difficulty becomes even more significant. ? (?) derived EB inequalities for $\beta$-mixing Hilbert-valued processes based on block techniques. Their bounds depend explicitly on the knowledge of mixing coefficients and the block length, and the effective sample size is reduced to the number of blocks, which in general prevents parametric (see Theorems 1-3 there).
@@ 1.3 Our Contributions
We develop simple self-normalized (empirical Bernstein) inequalities for both martingale difference and mixing dependence structures. The main contributions of our results are:
@@
    (weak dependence on support information) Our bounds are of the form
    $|\text{Empirical Mean}-\text{Marginal Mean}|\lesssim\sqrt{\frac{2\log(\frac{1}{\delta})\text{Var}_{n}}{n^{2}}}\lor\frac{2bt}{n},$
    where $\text{Var}_{n}$ is some “empirical variance” and $b$ is the uniform bound of random variables. xcept in regimes of extremely small variance or very small sample size, the leading term is governed entirely by the empirical variance. In particular, in many practical situations one does not need a precise numerical value of the support length $b$; the bound automatically ignores $b$ once $\text{Var}_{n}$ is in a typical range.
@@
    (No variance-proxy selection: empirical variance only) Our self-normalization avoids the variance–proxy selection problem that is central in existing empirical Bernstein and betting martingale approaches. The bound depends only on a simple empirical variance: for martingale difference sequences this is the usual sample variance, and for mixing sequences it is a block-based empirical long-run variance. We do not require any predictable variance proxy, nor any tuning of a variance process to ensure a super-martingale property.
@@
    (No rate assumptions of mixing coefficients) Our inequalities are valid without any explicit assumptions of the decay rate of mixing coefficients, like summability. At the same time, the self-normalizer adapts to the relevant variance scale: in the martingale (or i.i.d.) case it corresponds to the marginal variance, whereas in the mixing case it estimates the long-run variance via a simple block construction.
  • C5
    (Asymptotically Sharp) Under very mild conditions, the $\text{Var}_{n}$ introduced in C1 consistently estimates the relevant variance parameter, $\sigma^{*}$. i.e. The marginal variance in the martingale-difference/i.i.d. case and the long-run variance in the mixing case. As a consequence, for any $\delta>0$,
    $\sqrt{n}\left(\sqrt{\frac{2\log(\frac{1}{\delta})\text{Var}_{n}}{n^{2}}}\lor\frac{2bt}{n}\right)\xrightarrow[]{\mathbb{P}}\sigma^{*}\sqrt{2\log(\frac{1}{\delta})}.$
    Therefore, when the assumption of constant conditional expectation holds, our inequalities are asymptotically sharp in the sense of (5).
@@ 2.1 Martingale Difference
Given a probability space $(\Omega,\mathcal{F},\mathbb{P})$, let $(\mathcal{A}_{i})_{i\geq 0}$ be a filter satisfying $\mathcal{A}_{i}\subset\mathcal{F}$. Suppose $(Z_{i},\mathcal{A}_{i})_{i\geq 0}$ is a martingale difference sequence (MDS) such that $|Z_{i}|\leq b$ holds for any given $i\geq 1$. Then, by denoting $\mathbf{M}_{n}=\sum_{i=1}^{n}Z_{i}$, $\mathbf{M}]_{n}=\sum_{i=1}^{n}Z_{i}^{2}$, for any given $\delta>0$ and $n\geq 1$, we have
$\mathbb{P}\left(|\mathbf{M}_{n}|\leq\max\left\{\sqrt{2[\mathbf{M}]_{n}\log(\frac{1}{\delta})},2b\log(\frac{1}{\delta})\right\}\right)\geq 1-2\delta.$ (8)
As we can see in (8), combines the Bernstein–type tradeoff of Freedman’s inequality with a fully empirical self-normalization. It replaces the unobservable predictable quadratic variation $\langle\mathbf{M}\rangle_{n}$ by the quadratic variation $\mathbf{M}]_{n}$, which is completely determined by the data. Generally speaking, Theorem 1 will serve as the basic building block for our extensions to general mixing settings (Section 2.2), where $\mathbf{M}]_{n}$ is replaced by a block-based empirical long-run variance.
Of course, a more acceptable setting similar to MDS is that $\mathbb{E}[Z_{i}|\mathcal{A}_{i-1}]=\mu$ holds for every $i$, which makes $(Z_{i}-\mu,\mathcal{A}_{i})$ an MDS. However, when $\mu\neq 0$, $\mathbf{M}]_{n}$ becomes a term containing unknown parameter. Using empirical mean to replace $\mu$ yields the following empirical Bernstein inequality.
@@
Given a probability space $(\Omega,\mathcal{F},\mathbb{P})$, let $(\mathcal{A}_{i})_{i\geq 0}$ be a filter satisfying $\mathcal{A}_{i}\subset\mathcal{F}$. Suppose $(Z_{i})_{i\geq 0}$ is a process adapted to $(\mathcal{A}_{i})_{i\geq 0}$ and satisfies $\mathbb{E}[Z_{i}|\mathcal{A}_{i-1}]=\mu$ and $|Z_{i}|\leq b_{V}<\infty$. Then, by denoting $\overline{\mathbf{M}}_{n}=\frac{1}{n}\sum_{i=1}^{n}Z_{i}$ and $\mathbb{V}]_{n}=\sum_{i=1}^{n}(Z_{i}-\overline{\mathbf{M}}_{n})^{2}$, the following inequality holds for any given $t>0$ and $n\geq 1$,
@@
$\mathbb{P}\left(\Big|\overline{\mathbf{M}}_{n}-\mu\Big|\leq\nu_{n}(\delta)\max\left\{\sqrt{\frac{2[\mathbb{V}]_{n}\log(\frac{1}{\delta})}{n^{2}}},\frac{2b_{V}\log(\frac{1}{\delta})}{n}\right\}\right)\geq 1-2\delta,\ \ \nu_{n}(\delta)=\frac{1}{1-\sqrt{2\log(\frac{1}{\delta})/n}}.$ (9)
@@
It is instructive to compare Corollary 1 with Theorem 2 and Remark 1 of ? (?). In the bounded i.i.d. (or martingale-difference) case, their empirical Bernstein confidence sequence is asymptotically sharp in the sense of (5): for fixed $\delta$, the $\sqrt{n}$-scaled radius converges to $\sigma\sqrt{2\log(\frac{1}{\delta})}$. They work in a time-uniform setting and therefore construct a predictable variance proxy based on one-step-ahead predictions, together with a nontrivial betting/mixture strategy, and their target is the predictable (drifting) mean, which coincides with the marginal mean only under the martingale-difference assumption. Moreover, the sharpness discussion in Remark 1 is formulated under a homoscedasticity condition, where the conditional variance process converges to a constant so that the variance parameter is a single $\sigma^{2}$. The heteroscedastic case, with $\text{Var}(Z_{i})$ varying in $i$, is not treated explicitly. By contrast, Corollary 1 is purely fixed-$n$, focuses directly on the marginal mean, and uses only the naive empirical variance (plus a ”soft floor” depending on the range $b$, with no predictable variance proxy, no betting construction, and no tuning. For bounded independent or martingale-difference sequences with possibly non-constant marginal variances, the empirical variance automatically converges to the natural average scale $\frac{1}{n}\sum_{i=1}^{n}\text{Var}(Z_{i})=:\sigma^{2}_{n}$, and simultaneously, the $\sqrt{n}$-scaled radius converges to $\sigma_{n}\sqrt{2\log(\frac{1}{\delta})}$. In this sense, our empirical Bernstein bound is asymptotically sharp for general bounded independent or martingale-difference sequences, not only in the homoscedastic setting considered in ?’s sharpness discussion. Meanwhile, except for some extreme cases like significantly small variance and sample size, we can ignore the boundedness $b$ in high-probability bound, which is quite useful in the cases where we do not know the exact value of boundedness.
@@
However, many practitioners may be inclined to ignore the $\frac{2}{n}b_{V}\log(\frac{1}{\delta})$ directly. A natural question is the cost of this ignorance. Corollary 2 shows that, in a non-asymptotic level, the ignorance of $\frac{2}{n}b_{V}\log(\frac{1}{\delta})$ only cost an exponentially small loss of coverage.
@@
Suppose $(Z_{i})_{i=1}^{n}$ are independent copies of non-degenerated random variable $Z_{0}$ such that $|Z_{0}|\leq b$. By denoting $\text{Var}(Z_{0})=:\sigma^{2}$, $E|Z_{0}-\mu|^{4}=:m_{4}$,
$\mathbb{P}\left(\Big|\overline{\mathbf{M}}_{n}-\mu\Big|\leq\nu_{n}(\delta)\sqrt{\frac{2\log(\frac{1}{\delta})[\mathbb{V}]_{n}}{n^{2}}}\right)\geq 1-2\delta-\exp\left(-\frac{n(1-\eta)^{2}\sigma^{4}}{2(m_{4}+\frac{1}{3}b(1-\eta)\sigma^{2})}\right)$
holds for every $\delta>0$, $\eta\in(0,1)$ and $n\geq N(\delta):=\min\{n\geq 1:\eta\sigma^{2}\geq 2\frac{1}{n}\log(\frac{1}{\delta})b_{V}^{2}\}$, where $\nu_{n}(\delta)$ is introduced in Corollary 1.

### v3

# Self-Normalized Concentration Inequalities of Marginal Mean with Sample Variance Only

@@ Abstract
(This is the second version of this working paper. A more detailed follow-up with interesting applications is in preparation.) We develop a family of self-normalized concentration inequalities for marginal mean under martingale-difference structure and $\phi/\tilde{\phi}-$mixing conditions, where the latter includes many processes that are not strongly mixing. The variance term is fully data-observable: naive sample variance in the martingale case and an empirical block long-run variance under mixing conditions. Thus, no predictable variance proxy is required. No specific assumption on the decay of the mixing coefficients (e.g. summability) is needed for the validity. The constants are explicit and the bounds are ready to use.
@@ 1 Literature Review
This section is dedicated to the literature review of the previous research on self-normalized and empirical Bernstein (EB) inequalities for marginal mean of real-valued random variables.
@@ 1.1 Self-Normalized Inequality
They also proved that (3) holds for “heavy-on-left” MDS as well. Compared with (1), (2)-(4) hold under much weaker moment conditions but their statistical applications are not so convenient as (1). Meanwhile, the dependence on $\mathbf{M}]_{n}$ appears inside the Laplace transform or in the event $\{[\mathbf{M}]_{n}\leq y\}$, and there is no closed-form expression of the type $\sqrt{[\mathbf{M}]_{n}\log(\frac{1}{\delta})}$ with an explicit confidence parameter $\delta$. However, the variance ($\sigma$) and boundedness ($b$) are usually unknown to us, which reject the direct application of (1).
@@ 1.2 Empirical Bernstein Inequality of Marginal Mean
$\sqrt{n}L_{n}\xrightarrow[]{\mathbb{P}}\sigma\sqrt{2\log(1/\alpha)},$ (5)
@@
In summary, the results of Ramdas and co-authors are indeed elegant. But, as we discussed for fixed-n EB inequality of marginal mean, they have three structural limitations: (i) they require a known and tight bound on the support length to calculate the confidence interval; (ii) they primarily target the predictable mean (the arithmetic average of conditional expectations), and in general mixing settings there is no clear high-probability control of the gap between this predictable mean and the marginal mean—unless one assumes that all conditional expectations are constant, which essentially reduces to the martingale-difference case. Their recent important work, e.g. ? (?) and ? (?), still rely on the assumption of constant conditional expectation. (iii) they rely on a carefully tuned variance proxy satisfying the following three points simultaneously, (i) predictable; (ii) ensure the super-martingale property; (iii) converge to the true variance. The design of this proxy is highly non-trivial even for martingale differences or i.i.d. data. For some complex mixing processes, whose relevant variance of partial sum is the long-run variance rather than the marginal variance, the difficulty becomes even more significant.
As for EB inequality of dependent data, ? (?) derived EB inequalities for $\beta$-mixing Hilbert-valued processes based on block techniques. Their bounds depend explicitly on the knowledge of mixing coefficients and the block length, and the effective sample size is reduced to the number of blocks, which in general prevents parametric (see Theorems 1-3 there).
@@ 1.3 Our Contributions
The main contributions of this paper are:
@@
    (Two inequalities for MDS with clean constants) Our first contribution is a pair of clean self-normalized inequalities for martingale difference sequences. For all $t>0$ and $n\geq 1$,
    $\mathbb{P}\Big(\mathbf{M}_{n}\leq\sqrt{2\langle\mathbf{M}\rangle_{n}t}+\frac{b}{3}t\Big)\geq 1-2e^{-t}.$
    This inequality matches the classical Bernstein–Freedman constants while allowing the random predictable quadratic variation $\langle\mathbf{M}\rangle_{n}$ to appear directly in the bound. It provides a one-parameter form that does not follow from the traditional two-parameter or ratio-type self-normalized inequalities. We then upgrade this to a more useful empirical Bernstein inequality by replacing $\langle\mathbf{M}\rangle_{n}$ with $\mathbf{M}]_{n}$ (see (9)). The empirical version enlarges the constant only slightly, yet keeps the variance and rate dependence optimal, and crucially requires no variance proxy. These inequalities serve as the probabilistic backbone of all subsequent results, allowing us to derive empirical Bernstein bounds for much more complex dependent processes (see C2–C4).
@@
    (No variance-proxy selection: empirical variance only) Our EB inequality avoids the variance–proxy selection problem that is central in existing empirical Bernstein and betting martingale approaches. The bound depends only on a simple empirical variance: for martingale difference sequences this is the usual sample variance, and for mixing sequences it is a block-based empirical long-run variance. We do not require any predictable variance proxy, nor any tuning of a variance process to ensure a super-martingale property.
@@
    (No specific rate assumption of mixing coefficients) Our inequalities are valid without any explicit assumptions of the decay rate of mixing coefficients. At the same time, the self-normalizer adapts to the relevant variance scale under summability assumption: in the martingale (or i.i.d.) case it corresponds to the marginal variance, whereas in the mixing case it estimates the long-run variance via a simple block construction. Therefore, when the assumption of constant conditional expectation holds, our inequalities are asymptotically sharp in the sense of (5) up to a very mild constant.
@@ 2.1 Martingale Difference Sequence (MDS)
Given a probability space $(\Omega,\mathcal{F},\mathbb{P})$, let $(\mathcal{A}_{i})_{i\geq 0}$ be a filter satisfying $\mathcal{A}_{i}\subset\mathcal{F}$. Suppose $(Z_{i},\mathcal{A}_{i})_{i\geq 0}$ is a martingale difference sequence (MDS) such that $|Z_{i}|\leq b$ holds for any given $i\geq 1$. Then, based on the $\mathbf{M}_{n}=\sum_{i=1}^{n}Z_{i}$, $\langle\mathbf{M}\rangle_{n}=\sum_{i=1}^{n}\mathbb{E}[Z_{i}^{2}|\mathcal{A}_{i-1}]$ and $\mathbf{M}]_{n}=\sum_{i=1}^{n}Z^{2}_{i}$ introduced before in Section 1.1 (MDS implies zero mean), for any given $\delta>0$ and $n\geq 1$, we have
$\mathbb{P}\Big(|\mathbf{M}_{n}|\leq\sqrt{2\langle\mathbf{M}\rangle_{n}t}+\frac{bt}{3}\Big)\geq 1-2e^{-t},$ (8)
$\mathbb{P}\Big(|\mathbf{M}_{n}|\leq\sqrt{2[\mathbf{M}]_{n}t}+3.15bt\Big)\geq 1-3e^{-t}.$ (9)
It is still unclear whether the results in the previous research of self-normalized inequality could imply our (8) as a natural corollary without non-trivial work. Except for the literature mentioned in Section 1, ? (?) discuss self-normalized concentration phenomenon by only requiring $\mathbb{E}[\exp(\lambda A-\frac{1}{2}\lambda^{2}B^{2})]\leq 1$. But ? only concentrated on delivering moment or probabilistic inequalities of ratio-type statistics, which is hard to be extended to our settings directly. More recently, ? (?) talked about two parameter self-normalized inequalities. More specifically, by denoting $S_{n}(a)=c(a)\langle\mathbf{M}\rangle_{n}+[\mathbf{M}]_{n}$ ($c(a)$ a constant depending solely on $a>1/8$), ? gave a series exponential inequalities for ratio statistics (i.e. $\frac{\mathbf{M}_{n}}{S_{n}(a)}$) and two-parameter situations (e.g. ? (?). Hence, under martingale-difference assumption, it seems that our (8) is the first “one-parameter” (only a $t>0$ is needed) exponential inequality whose constants matche with classical Bernstein or Freedman inequality.
As we can see in (9), combines the Bernstein–type tradeoff of Freedman’s inequality with a fully empirical self-normalization. It replaces the unobservable predictable quadratic variation $\langle\mathbf{M}\rangle_{n}$ by the quadratic variation $\mathbf{M}]_{n}$, which is completely determined by the data. Generally speaking, Theorem 1 will serve as the basic building block for our extensions to general mixing settings (Section 2.2), where $\mathbf{M}]_{n}$ is replaced by a block-based empirical long-run variance. A more acceptable setting similar to MDS is that $\mathbb{E}[Z_{i}|\mathcal{A}_{i-1}]=\mu$ holds for every $i$, which makes $(Z_{i}-\mu,\mathcal{A}_{i})$ an MDS. However, when $\mu\neq 0$, $\mathbf{M}]_{n}$ becomes a term containing unknown parameter. Using empirical mean to replace $\mu$ yields the following empirical Bernstein inequality.
@@
Given a probability space $(\Omega,\mathcal{F},\mathbb{P})$, let $(\mathcal{A}_{i})_{i\geq 0}$ be a filter satisfying $\mathcal{A}_{i}\subset\mathcal{F}$. Suppose $(Z_{i})_{i\geq 0}$ is a process adapted to $(\mathcal{A}_{i})_{i\geq 0}$ and satisfies $\mathbb{E}[Z_{i}|\mathcal{A}_{i-1}]=\mu$ and $|Z_{i}|\leq b<\infty$. Then, by denoting $\overline{\mathbf{M}}_{n}=\frac{1}{n}\sum_{i=1}^{n}Z_{i}$ and $\mathbb{V}]_{n}=\sum_{i=1}^{n}(Z_{i}-\overline{\mathbf{M}}_{n})^{2}$, the following inequality holds for any given $t>0$ and $n\geq 1$, by denoting $\nu_{n}(\delta)=\frac{1}{1-\sqrt{2\log(\frac{1}{\delta})/n}}$, we have
@@
$\mathbb{P}\left(\Big|\overline{\mathbf{M}}_{n}-\mu\Big|\leq\nu_{n}(\delta)\left(\sqrt{\frac{2[\mathbb{V}]_{n}\log(\frac{1}{\delta})}{n^{2}}}+\frac{3.15b\log(\frac{1}{\delta})}{n}\right)\right)\geq 1-3\delta.$ (10)
@@
It is instructive to compare Corollary 1 with the Theorem 2, Remark 1 of ? (?). In the bounded i.i.d. (or martingale-difference) case, their empirical Bernstein confidence sequence is asymptotically sharp in the sense of (5): for fixed $\delta$, the $\sqrt{n}$-scaled radius converges to $\sigma\sqrt{2\log(\frac{1}{\delta})}$. They work in a time-uniform setting and therefore construct a predictable variance proxy based on one-step-ahead predictions, together with a nontrivial betting/mixture strategy, and their target is the predictable (drifting) mean, which coincides with the marginal mean only under the martingale-difference assumption. Moreover, the sharpness discussion in Remark 1 is formulated under a homoscedasticity condition, where the conditional variance process converges to a constant so that the variance parameter is a single $\sigma^{2}$. The heteroscedastic case, with $\text{Var}(Z_{i})$ varying in $i$, is not treated explicitly.
By contrast, Corollary 1 is purely fixed-$n$, focuses directly on the marginal mean, and uses only the naive empirical variance. Even though our constant is slightly enlarged (with probability no less than $1-2\alpha$, the constant part of the leading term is $\sigma\sqrt{2\log(1.5/\alpha)}$, which is slightly larger than (5)), we obtain a lot of important properties, like no predictable variance proxy, no betting construction, and no tuning. For bounded independent or martingale-difference sequences with possibly non-constant marginal variances, the empirical variance automatically converges to the natural average scale $\frac{1}{n}\sum_{i=1}^{n}\text{Var}(Z_{i})=:\sigma^{2}_{n}$. In this sense, our empirical Bernstein bound is asymptotically sharp up to a minor constant ($\sqrt{\log(\frac{1.5}{\alpha})/\log(\frac{1}{\alpha})}$) for general bounded independent or martingale-difference sequences, not only in the homoscedastic setting.
@@
However, many practitioners may be inclined to ignore the $\frac{3.15b}{n}\log(\frac{1}{\delta})$ directly. A natural question is the cost of this ignorance. Corollary 2 shows that, in a non-asymptotic level, the ignorance of $\frac{3.15b}{n}\log(\frac{1}{\delta})$ only cost an exponentially small loss of coverage.
@@
Suppose $(Z_{i})_{i=1}^{n}$ are independent copies of non-degenerated random variable $Z_{0}$ such that $|Z_{0}|\leq b$. By denoting $\text{Var}(Z_{0})=:\sigma^{2}$, $E|Z_{0}-\mu|^{4}=:m_{4}$, for every user-defined $\delta>0$, $\eta\in(0,1)$, $\xi_{n}=o(\frac{1}{\sqrt{n}})$,
$\mathbb{P}\left(\Big|\overline{\mathbf{M}}_{n}-\mu\Big|\leq\nu_{n}(\delta)(1+\xi_{n})\sqrt{\frac{2\log(\frac{1}{\delta})[\mathbb{V}]_{n}}{n^{2}}}\right)\geq 1-3\delta-\exp\left(-\frac{n(1-\eta)^{2}\sigma^{4}}{2(m_{4}+\frac{1}{3}b(1-\eta)\sigma^{2})}\right)$
holds when $n\geq N(\delta,\eta)=\min\{n\geq 1:\eta\text{Var}(Z_{0})\geq\frac{5b^{2}\log(\frac{1}{\delta})}{n\xi_{n}^{2}}\}$. $\nu_{n}(\delta)$ is introduced in Corollary 1.